[figure: 9.7 cm]

THE ULTRA WEAK REMANENT MAGNETIC MOMENT IN QUASI-2D MAGNETOELECTRIC CRYSTAL LiNiP O_{4}

Yu. Kharchenko^{a}, N.F. Kharchenko^{a} and M. Baran^{b}, R. Szymczak^{b} aInst. of Low Temp. Phys. & Eng., NASU, Pr.Lenin 47, 61103 Kharkiv, Ukraine bInst. of Physics, Polish Acad. of Sci., Al.Lotnikow 32/46, 02-668 Warsaw, Poland

Magnetization of the single crystal of LiNiP O_{4} as a function of a temperature and a magnetic field has been measured along the crystallographic axes using the SQUIDmagnetometer. The ultra weak remanent magnetic moment in LiNiP O_{4} has been revealed. The measured residual magnetization was greatest along the c axis. Its value is about 0.005 G at 5 K. Along the crystallographic a and b axes its magnitueds were less on order than along axis c. With increase of a temperature of the sample up to the Neel temperature a magnitude of the remanent magnetization was decreasing monotonously to zero. The nature of its occurrence is not clear. The group of magnetic symmetry P nm′a that was established for this crystal at neutron-diffraction studies forbids weak ferromagnetism. The temperature behaviour of observed remanent magnetization essentially differs from that which was observed in the lightly iron-doped crystals. It suggests that the observed effect is not caused by uncontrollable impurities. It is possible, that the occurrence of the residual moment is connected with weak modulation of a spin structure of the crystal. Let’s note that the incommensurate structure in LiNiP O_{4} was revealed recently in a narrow temperature range close to T_{N} at the neutron-diffraction studies. The presence of weak ferromagnetic moment in this crystal gives hope for explanations of abnormal behaviour of the magnetoelectric effect in a magnetic field observed in this crystal.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : postar

Corresponding author : Yu.Kharchenko

Address for correspondence : Institute for Low Temperature Physics and Engineering of the Academy of Sciences of Ukraine Pr. Lenin 47, 61103 Kharkiv Ukraine

Email address : ykharchenko@ilt.kharkov.ua